OpenStudy (calculusxy):

If the population of Country X increased by 10 percent each year over a 2-year period, what was the total percent increase in the population over the entire period? @shamil98

OpenStudy (anonymous):

21% Example. 100x10%...10. 100+10=110 110x10%...11. 110+11=121 121-100=21 21/100=21%